Each firm in a competitive market has a cost


function of C = q + q2 + q3. The market has an


unlimited number of potential firms. The market


demand function is Q = 24 - p. Determine the


long-run equilibrium price, quantity per firm, market quantity, and number of firms. How do these


values change if a tax of $1 per unit is collected from


each firm? (Hint: See Solved Problem 8.4.) M

1
Expert's answer
2021-10-31T18:29:05-0400

Solution:

In the long-run equilibrium price, the profits are zero, therefore: P = MC = ATC


MC = "\\frac{\\partial TC} {\\partial Q} =\n\n\u200b"1 + 2q + 3q2

ATC = 1 + q + q2


Set: MC = ATC

1 + 2q + 3q2 = 1 + q + q2

q = 0.5

Q = 0.5

The long run equilibrium output per firm = 0.5 units

The long run equilibrium price is = MC = 1 + 2q + 3q2 = 1 + 2(0.5) + 3(0.52) = 2.75

The long run equilibrium price = 2.75

 

The market quantity is determined through the market demand, Q = 24 – p

Q = 24 – 2.75 = 21.25

The number of firms = "\frac{21.25}{0.5} = 43" firms